Colter Steel has $4,900,000 in assets. Temporary current assets Permanent current assets Fixed assets $ 1,800,000 1,540,000 1,560,000 $ 4,900,000 Total assets Assume the term structure of interest rates becomes inverted, with short-term rates going to 14 percent and long-term rates 4 percentage points lower than short-term rates. Earnings before interest and taxes are $1,040,000. The tax rate is 30 percent If long-term financing is perfectly matched (synchronized) with long-term asset needs, and the same is true of short-term financing, what will earnings after taxes be? Earnings after taxes
Answer #1

Total short term liabilities = $1800,000

Interest on short term liabilities = $1800,000*14% = $252,000

Total Long term liabilities (1,540,000+1,560,000)= $3,100,000

Interest on Total Long term liabilities = $3,100,000*10% =$310,000

Total interest expense = $310,000 + $252,000 = $562,000

Earnings before interest and taxes = $1,040,000

:Less: Interest Expense = $562,000

Earnings before tax = $478,000

Less Income Tax(30%) = $143,400

Earnings after taxes = $334,600